### v1

@@ 5.1 Current-free boundary conditions
Here, we outline the approach adopted in the 3D code MATINS (Dehman et al, 2023c). As the external potential solution is, by definition, both solenoidal and irrotational, the magnetic field can be expressed as the gradient of a magneto-static potential $\chi$ which obeys the Laplace equation (see Appendix B for more details). One can then expand the scalar function $\chi$ in spherical harmonics as follows:
@@
so that, we can evaluate the coefficients by applying the orthogonality properties of spherical harmonics to Eq. 53. Integrating over the star surface one can obtain:
@@
$(1-z)\frac{\partial^{2}\Phi_{\ell m}}{\partial r^{2}}+\frac{z}{r}\frac{\partial\Phi_{\ell m}}{\partial r}-\frac{\ell(\ell+1)}{r^{2}}\Phi_{\ell m}=0~,$ (57)
where we assume the metric (1), and $z\equiv\frac{2GM}{c^{2}r}$. In this subsection, we explicitly reintroduce the relativistic corrections, as they will play an important role in the spin-down rate, discussed later in 5.4.1.
@@
$\Phi_{1}$ $=$ $C_{1}r^{2}\left[\ln(1-z)+z+\frac{z^{2}}{2}\right]~,$ (59)
@@
$\left.\frac{\partial\Phi_{\ell}}{\partial r}\right|_{r=R}=-\frac{\ell}{R}f_{\ell}\Phi_{\ell}~,$ (63)
where the $f_{\ell}$’s are relativistic corrections that only depend on the value of $z$ at the star surface, $z(r=R)$ (in the Newtonian limit all $f_{\ell}=1$), and can be evaluated numerically only once with the help of any algebraic manipulator and stored [^9] See Rädler et al 2001 for an alternative form to evaluate $f_{\ell}$ based on the expansion in a series of powers of $1/r$..
5.2 Force–free boundary conditions
In axial symmetry, the construction of force-free (FF) magnetospheres for (non-rotating) magnetars is a well-studied problem, even in the relativistic case (see, e.g., Kojima 2017 and references therein). In the context of magneto-thermal evolution, Akgün et al (2018) explored a method to impose such boundary conditions by solving the Grad-Shafranov equation, at each time step, to match the internal evolution of the star. Let us review their approach. Considering axial symmetry, the magnetic field can be written as follows:
$\vec{B}=\frac{(\partial P/\partial\theta)}{r^{2}\sin\theta}{\hat{r}}-\frac{(\partial P/\partial r)}{r\sin\theta}{\hat{\theta}}+\frac{T}{r\sin\theta}{\hat{\varphi}}~$, (64)
@@ 5.2 Force–free boundary conditions
In principle, there is an infinite family of external force-free solutions for a given radial magnetic field at the surface, because of the freedom to choose the functional form of $T(P)$. The main problem of this approach is how to continuously match the arbitrary field configuration, resulting from the evolution in the crust, while enforcing the force-free solution outside. In the crust, any line bundle marked by a given magnetic flux $P$ has in general different values of $T$ because, internally, the force-free condition does not hold. As discussed in Akgün et al (2018), there is an intrinsic inconsistency in the possibly multi-valued function $T(P)$, if we strictly take it from the values at the surface ($r=R$). They address this problem by symmetrizing the numerical function $T(P)$, which is physically equivalent to allowing the propagation through the surface only of the modes compatible with solutions of the Grad-Shafranov equation.
@@
In Fig. 13, we show a representative result, showing the evolution of an axisymmetric magnetospheric configuration physically connected to the interior. The initial model consists of both poloidal and toroidal dipolar components, with the latter extending beyond the surface. As the internal magnetic field evolves, the external magnetic field is consistently twisted, by the injection of magnetic helicity (i.e., currents) in the magnetosphere. Solutions are calculated at each time step until a critical point, where numerical solutions cannot be found anymore. The absence of a compatible solution physically means that the magnetosphere is expected to become unstable, possibly resulting in a global reconfiguration by opening of the twisted field lines and magnetic reconnection. Such reconfiguration, occurring on dynamical timescales (ms), is of extreme interest for the observed transient phenomenology of magnetars (Rea and Esposito, 2011), but cannot be simulated with long-term evolution codes. These processes have been studied in detail in dedicated force-free electrodynamics simulations, in both the Newtonian and general relativistic (GR) cases (Parfrey et al, 2013a; Carrasco et al, 2019).
@@
Figure 13: Evolution of a twisted magnetosphere with coupling with the interior. The left and right panels show snapshots at $t=0$ and $t\sim 1.58$ kyr, the critical time when the magnetosphere of this particular model has stored the maximum possible twist. Figure courtesy of Akgün et al (2017).
@@
Stefanou et al (2023a) presented a comprehensive study of force-free twisted magnetar magnetospheres with non-linear current distributions. Their work solves the force-free equations within a compactified spherical coordinate system, using the Grad–Rubin method. At the stellar surface, they applied suitable boundary conditions to prescribe both the current distribution and the magnetic field. The method’s accuracy is verified by reproducing a range of established analytical solutions and axisymmetric numerical results. Building on this validation, they explore fully 3D configurations with non-axisymmetric current patterns—for example, magnetic fields with localized twists that mimic surface hotspots. The study analyzes key physical quantities, including magnetic energy, helicity, and twist, and considers the implications for the magnetar’s energy budget, surface heating, and magnetic diffusion timescales, all in connection with possible observational signatures.
@@
In a recent paper (Stefanou et al, 2025), building on the methodology of previous works in axisymmetry (Stefanou et al, 2023c; Urbán et al, 2023), the authors employ a novel methodology based on PINNs to model pulsar and magnetar magnetospheres, spanning both axisymmetric and fully three-dimensional configurations for stars of varying compactness. The force-free equations are directly solved in the form
@@
These works highlight the promise of PINNs as an efficient and generalizable tool for simulating 3D magnetospheres (or other elliptic PDE problems), offering new future perspectives to investigate the magnetar phenomenology. The preliminary (axisymmetric) results presented in Urbán et al (2023) demonstrate that this approach can be applied to the astrophysical problem of magnetic field evolution within a NS interior, coupled to a force-free magnetosphere. Using a PINN reduced the computational cost by more than an order of magnitude compared to a finite difference scheme applied to a similar case. These findings open the door to future 3D extensions of this or related problems, where implementing generalized boundary conditions is otherwise prohibitively expensive.
@@ 5.3 Dynamic force-free relaxation
The idea is that, since the magnetospheric timescales are many orders of magnitude shorter than the interior, the long-term evolution of the magnetosphere can be seen as a series of equilibrium states, attained immediately after every time step of the interior. Therefore, one can activate an artificial term that dynamically leads to the force-free solution. This approach is similar to the magneto-frictional method (Yang et al, 1986; Roumeliotis et al, 1994), as known in solar physics. The modified induction equation employed in the exterior of the star has a mathematical structure equivalent to an ambipolar term (as in Eq. (29)), which forces currents to gradually align to magnetic field lines, without having to solve the elliptical Grad-Shafranov equation at every time step (which is numerically expensive). This also allows us to account for the transfer of helicity and provides a mechanism to continuously feed currents that twist the magnetosphere. The caveat is that the coupling coefficient (quantifying the ratio between the interior and exterior timescales) must be fine-tuned to prevent the exterior dynamics from being neither too fast (it would excessively limit the time step), nor too slow (it would not manage to relax to a force-free configuration and would cause non-negligible, unphysical feedback on the interior). In the NS long-term evolution scenario, such a strategy has only been explored preliminarily in the 3D Cartesian parallelized code used in Viganò et al (2019). More studies are needed to test the feasibility of this approach.
@@ 5.4 Evolution of spin period and obliquity
The magnetic field evolution also influences the rotational properties of the star. As NSs age, they lose angular momentum through electromagnetic torques, i.e. they spin down (Spitkovsky, 2006; Beskin et al, 2013; Philippov et al, 2014). The poloidal dipolar magnetic field dominates this process, since higher-order multipoles decay rapidly with distance. Since the period and its derivative are among the primary observables, it is relevant to examine the standard approximations and their limitations. The basic argument for quantitatively estimating the dipolar component of the surface field at the magnetic pole, $B_{p}$, involves equating the rotational energy losses, $I\Omega\dot{\Omega}$ (where $I$ is the moment of inertia of the interior component coupled to the magnetosphere, $\Omega$ is the spin angular velocity, and $\dot{\Omega}$ its time derivative), to the electromagnetic torque, which is
$L_{sd}=\frac{B_{p}^{2}R^{6}\Omega^{4}}{6c^{3}}f_{\phi}~,$ (67)
@@
A rotating plasma-filled magnetosphere has no trivial solution; therefore, numerical simulations are necessary. In a more general and realistic case, $f_{\phi}=\kappa_{0}+\kappa_{1}\sin^{2}\phi$ (Spitkovsky, 2006), and the coupled evolution of spin period and inclination angle is governed by (Philippov et al, 2014):
$\dot{P}=\beta\frac{B_{p}^{2}}{P}(\kappa_{0}+\kappa_{1}\sin^{2}{\phi})~,$ (69)
$\dot{\chi}=-\kappa_{2}\beta\frac{B_{p}^{2}}{P^{2}}\sin{\phi}\cos{\phi}~,$ (70)
@@
$\beta\equiv\frac{\pi^{2}R^{6}}{Ic^{3}}~,$ (71)
and the coefficients $\kappa_{0}$, $\kappa_{1}$, $\kappa_{2}$ depend on the magnetosphere geometry and its physical conditions and determine the magnetospheric torque. In the classical (unphysical) vacuum dipole model $\kappa_{0}=0$, $\kappa_{1}=\kappa_{2}=2/3$, which incorrectly suggests that an aligned rotator ($\phi=0$) experiences no torque and would not spin down (e.g. Johnston and Karastergiou, 2017, but similar arguments are abundant). This is physically inaccurate, as realistic 3D numerical models of plasma-filled magnetospheres find $\kappa_{0}\sim\kappa_{1}\approx 1$ (Spitkovsky, 2006; Philippov et al, 2014), with $\kappa_{2}$ ranging from 0 to 1. Many groups achieve comparable results despite using different methods (force-free electrodynamics vs. particle-in-cell simulations) and physical components (resistive or purely force-free, relativistic or non-relativistic).
Moreover, Philippov et al (2014) demonstrated that the alignment of the rotation and magnetic axes in a vacuum magnetosphere model occurs much faster (exponentially, with characteristic time $\tau_{0}=\frac{P_{0}}{2\beta B_{0}}$) compared to realistic plasma-filled magnetospheres (where the alignment angle decreases following a power-law). In a realistic case, variations in the inclination angle typically cause torque corrections of up to a factor of $\approx 2$ (similar to the uncertainty in the NS moment of inertia). Alignment cannot halt the star’s period evolution.
@@ 5.4.1 General Relativistic effects
Several advanced GR force-free electrodynamics or particle-in-cell magnetospheric simulations (e.g., Ruiz et al 2014; Philippov et al 2015; Carrasco et al 2018) indicate that the spin-down luminosity, measured through the Poynting flux at the light cylinder, exceeds that of Newtonian models, showing higher values of $\kappa_{0}$ and $\kappa_{1}$ (see Table 2 in Pétri, 2016). The physical reason for this is the larger fraction of open field lines as the compactness ratio $M/R$ increases, with $M$ being the mass of the star and $R$ the areal radius. Care must be taken when interpreting quantities that depend on the reference frame. The relativistic magnetospheric simulations mentioned above yield results expressed as the magnetic moment observed at infinity. Consequently, the reported slight increase in $L_{sd}$ with compactness must be interpreted in terms of the observable quantities in a specific frame.
@@
$N_{R}:=\frac{\Omega_{0}}{\Omega_{*}}=\sqrt{1-z}<1~,$ (73)
@@
Using representative parameters ($\phi=\pi/4$, $M=1.4~M_{\odot}$, $R=12$ km, $I=0.4~MR^{2}$, e.g. Lattimer and Prakash 2001; Bejger and Haensel 2002), Stefanou et al (2025) find that a GR-corrected inferred $B_{p}$ reads
@@
Note that these GR corrections affect the entire NS population, potentially introducing significant bias in the inferred magnetic fields. Additional effects may further skew the inferred fields in magnetars. On one hand, in a highly twisted magnetosphere, the Poynting flux can be further amplified as toroidal pressures expand field lines beyond the light cylinder (Parfrey et al, 2013b). For instance, Ntotsikas and Gourgouliatos (2025) found that extreme twists could increase the spin-down luminosity by a factor of up to 16, which would further amplify this correction in some cases. On the other hand, particle winds produce a similar effect, particularly pronounced in magnetars (Tong et al, 2013). The combined impact of all effects likely results in a significant systematic overestimation of magnetic field strength when applying the classical Newtonian dipole model (Pétri, 2019).

### v2

@@ 5.1 Current-free boundary conditions
Here, we outline the approach adopted in the 3D code MATINS (Dehman et al., 2023c). As the external potential solution is, by definition, both solenoidal and irrotational, the magnetic field can be expressed as the gradient of a magneto-static potential $\chi$ which obeys the Laplace equation (see Appendix B for more details). One can then expand the scalar function $\chi$ in spherical harmonics as follows:
@@
so that, we can evaluate the coefficients by applying the orthogonality properties of spherical harmonics to Eq. (53). Integrating over the star surface one can obtain:
@@
$(1-z)\frac{\partial^{2}\Phi_{\ell m}}{\partial r^{2}}+\frac{z}{r}\frac{\partial\Phi_{\ell m}}{\partial r}-\frac{\ell(\ell+1)}{r^{2}}\Phi_{\ell m}=0\,$, (57)
where we assume the metric (1), and $z\equiv\frac{2GM}{c^{2}r}$. In this subsection, we explicitly reintroduce the relativistic corrections, as they will play an important role in the spin-down rate, discussed later in Sect. 5.4.1.
@@
$\Phi_{1}$ $=$ $C_{1}r^{2}\left[\ln(1-z)+z+\frac{z^{2}}{2}\right]\,$, (59)
@@
$\left.\frac{\partial\Phi_{\ell}}{\partial r}\right|_{r=R}=-\frac{\ell}{R}f_{\ell}\Phi_{\ell}\,$, (63)
where the $f_{\ell}$’s are relativistic corrections that only depend on the value of $z$ at the star surface, $z(r=R)$ (in the Newtonian limit all $f_{\ell}=1$), and can be evaluated numerically only once with the help of any algebraic manipulator and stored [^10] See Rädler et al. 2001 for an alternative form to evaluate $f_{\ell}$ based on the expansion in a series of powers of $1/r$..
5.2 Force-free boundary conditions
In axial symmetry, the construction of force-free (FF) magnetospheres for (non-rotating) magnetars is a well-studied problem, even in the relativistic case (see, e.g., Kojima 2017 and references therein). In the context of magneto-thermal evolution, Akgün et al. (2018) explored a method to impose such boundary conditions by solving the Grad-Shafranov equation, at each time step, to match the internal evolution of the star. Let us review their approach. Considering axial symmetry, the magnetic field can be written as follows:
$\vec{B}=\frac{(\partial P/\partial\theta)}{r^{2}\sin\theta}{\hat{r}}-\frac{(\partial P/\partial r)}{r\sin\theta}{\hat{\theta}}+\frac{T}{r\sin\theta}{\hat{\varphi}}\,$, (64)
@@ 5.2 Force-free boundary conditions
In principle, there is an infinite family of external force-free solutions for a given radial magnetic field at the surface, because of the freedom to choose the functional form of $T(P)$. The main problem of this approach is how to continuously match the arbitrary field configuration, resulting from the evolution in the crust, while enforcing the force-free solution outside. In the crust, any line bundle marked by a given magnetic flux $P$ has in general different values of $T$ because, internally, the force-free condition does not hold. As discussed in Akgün et al. (2018), there is an intrinsic inconsistency in the possibly multi-valued function $T(P)$, if we strictly take it from the values at the surface ($r=R$). They address this problem by symmetrizing the numerical function $T(P)$, which is physically equivalent to allowing the propagation through the surface only of the modes compatible with solutions of the Grad-Shafranov equation.
@@
In Fig. 13, we show a representative result, showing the evolution of an axisymmetric magnetospheric configuration physically connected to the interior. The initial model consists of both poloidal and toroidal dipolar components, with the latter extending beyond the surface. As the internal magnetic field evolves, the external magnetic field is consistently twisted, by the injection of magnetic helicity (i.e., currents) in the magnetosphere. Solutions are calculated at each time step until a critical point, where numerical solutions cannot be found anymore. The absence of a compatible solution physically means that the magnetosphere is expected to become unstable, possibly resulting in a global reconfiguration by opening of the twisted field lines and magnetic reconnection. Such reconfiguration, occurring on dynamical timescales (ms), is of extreme interest for the observed transient phenomenology of magnetars (Rea and Esposito, 2011), but cannot be simulated with long-term evolution codes. These processes have been studied in detail in dedicated force-free electrodynamics simulations, in both the Newtonian and general relativistic (GR) cases (Parfrey et al., 2013; Carrasco et al., 2019).
@@
Figure 13: Evolution of a twisted magnetosphere with coupling with the interior. The left and right panels show snapshots at $t=0$ and $t\sim 1.58$ kyr, the critical time when the magnetosphere of this particular model has stored the maximum possible twist. Image reproduced with permission from Akgün et al. (2017), copyright by the author(s).
@@
Stefanou et al. (2023a) presented a comprehensive study of force-free twisted magnetar magnetospheres with non-linear current distributions. Their work solves the force-free equations within a compactified spherical coordinate system, using the Grad–Rubin method. At the stellar surface, they applied suitable boundary conditions to prescribe both the current distribution and the magnetic field. The method’s accuracy is verified by reproducing a range of established analytical solutions and axisymmetric numerical results. Building on this validation, they explore fully 3D configurations with non-axisymmetric current patterns—for example, magnetic fields with localized twists that mimic surface hotspots. The study analyzes key physical quantities, including magnetic energy, helicity, and twist, and considers the implications for the magnetar’s energy budget, surface heating, and magnetic diffusion timescales, all in connection with possible observational signatures.
@@
In a recent paper (Stefanou et al., 2025), building on the methodology of previous works in axisymmetry (Stefanou et al., 2023b; Urbán et al., 2023), the authors employ a novel methodology based on PINNs to model pulsar and magnetar magnetospheres, spanning both axisymmetric and fully three-dimensional configurations for stars of varying compactness. The force-free equations are directly solved in the form
@@
These works highlight the promise of PINNs as an efficient and generalizable tool for simulating 3D magnetospheres (or other elliptic PDE problems), offering new future perspectives to investigate the magnetar phenomenology. The preliminary (axisymmetric) results presented in Urbán et al. (2023) demonstrate that this approach can be applied to the astrophysical problem of magnetic field evolution within a NS interior, coupled to a force-free magnetosphere. Using a PINN reduced the computational cost by more than an order of magnitude compared to a finite difference scheme applied to a similar case. These findings open the door to future 3D extensions of this or related problems, where implementing generalized boundary conditions is otherwise prohibitively expensive.
@@ 5.3 Dynamic force-free relaxation
The idea is that, since the magnetospheric timescales are many orders of magnitude shorter than the interior, the long-term evolution of the magnetosphere can be seen as a series of equilibrium states, attained immediately after every time step of the interior. Therefore, one can activate an artificial term that dynamically leads to the force-free solution. This approach is similar to the magneto-frictional method (Yang et al., 1986; Roumeliotis et al., 1994), as known in solar physics. The modified induction equation employed in the exterior of the star has a mathematical structure equivalent to an ambipolar term (as in Eq. (29)), which forces currents to gradually align to magnetic field lines, without having to solve the elliptical Grad-Shafranov equation at every time step (which is numerically expensive). This also allows us to account for the transfer of helicity and provides a mechanism to continuously feed currents that twist the magnetosphere. The caveat is that the coupling coefficient (quantifying the ratio between the interior and exterior timescales) must be fine-tuned to prevent the exterior dynamics from being neither too fast (it would excessively limit the time step), nor too slow (it would not manage to relax to a force-free configuration and would cause non-negligible, unphysical feedback on the interior). In the NS long-term evolution scenario, such a strategy has only been explored preliminarily in the 3D Cartesian parallelized code used in Viganò et al. (2019). More studies are needed to test the feasibility of this approach.
@@ 5.4 Evolution of spin period and obliquity
The magnetic field evolution also influences the rotational properties of the star. As NSs age, they lose angular momentum through electromagnetic torques, i.e. they spin down (Spitkovsky, 2006; Beskin et al., 2013; Philippov et al., 2014). The poloidal dipolar magnetic field dominates this process, since higher-order multipoles decay rapidly with distance. Since the period and its derivative are among the primary observables, it is relevant to examine the standard approximations and their limitations. The basic argument for quantitatively estimating the dipolar component of the surface field at the magnetic pole, $B_{p}$, involves equating the rotational energy losses, $I\Omega\dot{\Omega}$ (where $I$ is the moment of inertia of the interior component coupled to the magnetosphere, $\Omega$ is the spin angular velocity, and $\dot{\Omega}$ its time derivative), to the electromagnetic torque, which is
$L_{sd}=\frac{B_{p}^{2}R^{6}\Omega^{4}}{6c^{3}}f_{\phi}\,$, (67)
@@
A rotating plasma-filled magnetosphere has no trivial solution; therefore, numerical simulations are necessary. In a more general and realistic case, $f_{\phi}=\kappa_{0}+\kappa_{1}\sin^{2}\phi$ (Spitkovsky, 2006), and the coupled evolution of spin period and inclination angle is governed by (Philippov et al., 2014):
$\dot{P}=\beta\frac{B_{p}^{2}}{P}(\kappa_{0}+\kappa_{1}\sin^{2}{\phi})\,$, (69)
$\dot{\chi}=-\kappa_{2}\beta\frac{B_{p}^{2}}{P^{2}}\sin{\phi}\cos{\phi}\,$, (70)
@@
$\beta\equiv\frac{\pi^{2}R^{6}}{Ic^{3}}\,$, (71)
and the coefficients $\kappa_{0}$, $\kappa_{1}$, $\kappa_{2}$ depend on the magnetosphere geometry and its physical conditions and determine the magnetospheric torque. In the classical (unphysical) vacuum dipole model $\kappa_{0}=0$, $\kappa_{1}=\kappa_{2}=2/3$, which incorrectly suggests that an aligned rotator ($\phi=0$) experiences no torque and would not spin down (e.g. Johnston and Karastergiou, 2017, but similar arguments are abundant). This is physically inaccurate, as realistic 3D numerical models of plasma-filled magnetospheres find $\kappa_{0}\sim\kappa_{1}\approx 1$ (Spitkovsky, 2006; Philippov et al., 2014), with $\kappa_{2}$ ranging from 0 to 1. Many groups achieve comparable results despite using different methods (force-free electrodynamics vs. particle-in-cell simulations) and physical components (resistive or purely force-free, relativistic or non-relativistic).
Moreover, Philippov et al. (2014) demonstrated that the alignment of the rotation and magnetic axes in a vacuum magnetosphere model occurs much faster (exponentially, with characteristic time $\tau_{0}=\frac{P_{0}}{2\beta B_{0}}$) compared to realistic plasma-filled magnetospheres (where the alignment angle decreases following a power-law). In a realistic case, variations in the inclination angle typically cause torque corrections of up to a factor of $\approx 2$ (similar to the uncertainty in the NS moment of inertia). Alignment cannot halt the star’s period evolution.
@@ 5.4.1 General Relativistic effects
Several advanced GR force-free electrodynamics or particle-in-cell magnetospheric simulations (e.g., Ruiz et al. 2014; Philippov et al. 2015; Carrasco et al. 2018) indicate that the spin-down luminosity, measured through the Poynting flux at the light cylinder, exceeds that of Newtonian models, showing higher values of $\kappa_{0}$ and $\kappa_{1}$ (see Table 2 in Pétri, 2016). The physical reason for this is the larger fraction of open field lines as the compactness ratio $M/R$ increases, with $M$ being the mass of the star and $R$ the areal radius. Care must be taken when interpreting quantities that depend on the reference frame. The relativistic magnetospheric simulations mentioned above yield results expressed as the magnetic moment observed at infinity. Consequently, the reported slight increase in $L_{sd}$ with compactness must be interpreted in terms of the observable quantities in a specific frame.
@@
$N_{R}:=\frac{\Omega_{0}}{\Omega_{*}}=\sqrt{1-z}<1\,$, (73)
@@
Using representative parameters ($\phi=\pi/4$, $M=1.4~M_{\odot}$, $R=12$ km, $I=0.4~MR^{2}$, e.g. Lattimer and Prakash 2001; Bejger and Haensel 2002), Stefanou et al. (2025) find that a GR-corrected inferred $B_{p}$ reads
@@
Note that these GR corrections affect the entire NS population, potentially introducing significant bias in the inferred magnetic fields. Additional effects may further skew the inferred fields in magnetars. On one hand, in a highly twisted magnetosphere, the Poynting flux can be further amplified as toroidal pressures expand field lines beyond the light cylinder (Parfrey et al., 2013). For instance, Ntotsikas and Gourgouliatos (2025) found that extreme twists could increase the spin-down luminosity by a factor of up to 16, which would further amplify this correction in some cases. On the other hand, particle winds produce a similar effect, particularly pronounced in magnetars (Tong et al., 2013). The combined impact of all effects likely results in a significant systematic overestimation of magnetic field strength when applying the classical Newtonian dipole model (Pétri, 2019).